### v2

# MathOPEval: A Fine-grained Evaluation Benchmark for Visual Operations of MLLMs in Mathematical Reasoning

Abstract
Recent progress in Multi-modal Large Language Models (MLLMs) has enabled step-by-step multi-modal mathematical reasoning by performing visual operations based on the textual instructions. A promising approach uses code as an intermediate representation to precisely express and manipulate the images in the reasoning steps. However, existing evaluations focus mainly on text-only reasoning outputs, leaving the MLLM’s ability to perform accurate visual operations via code largely unexplored. This work takes a first step toward addressing that gap by evaluating MLLMs’ code-based capabilities in multi-modal mathematical reasoning. Specifically, our framework focuses on two key evaluation aspects: (1) Multi-modal Code Generation (MCG) evaluates the model’s ability to accurately understand and construct visualizations from scratch. (2) Multi-modal Code Editing (MCE) assesses the model’s capacity for fine-grained operations, which include three types: Deletion, Modification and Annotation. To evaluate the above tasks, we incorporate a dataset that covers the five most popular types of mathematical figures, including geometric diagrams, function plots, and three types of statistical charts, to provide a comprehensive and effective measurement of existing MLLMs. Our experimental evaluation involves nine mainstream MLLMs, and the results reveal that existing models still lag significantly behind human performance in performing fine-grained visual operations.
publicationid: pubid: 0000–0000/00\$00.00 © 2021 IEEE
I Introduction
RECENT years have witnessed remarkable progress in visual mathematical reasoning by Multi-modal Large Language Models (MLLMs) [1, 2, 3, 4]. Given an image and a textual problem description, MLLMs generate step-by-step reasoning toward the solution. These reasoning steps often include visual operations such as annotating angles, drawing auxiliary lines, and identifying key elements (Fig. 1). Compared to text-only mathematical reasoning, this task demands stronger multi-modal capabilities, particularly in aligning textual and visual information, planning and conducting visual operations to reach the correct answer.
Figure 1: Comparison of the paradigm of (a) multi-modal input, text-only output and (b) multi-modal input, multi-modal output with four types of visual operations.
Recently, using code as an intermediate representation for multi-modal mathematical reasoning has emerged as a promising approach, leading to substantial performance gains [5, 6]. In this paradigm, images are translated into corresponding Python or  codes that can accurately reconstruct them. Visual operations are then represented as code edits. Code serves as a precise and structured textual representation of images, aligning well with language-based reasoning and reducing the ambiguity often present in direct visual operations. However, recent evaluations of multi-modal mathematical reasoning have primarily focused on text-only outputs[5, 7, 8, 6]. It remains unclear how effectively current MLLMs can generate intermediate code that reflects accurate visual operations—an essential step toward precise, interpretable multi-modal mathematical reasoning (Fig. 1).
Figure 2: Illustration of initial dataset for four visual operations across five visualization types. The dataset includes evaluation instructions, code, and images constructed for different tasks and visualization types. Please refer to Section III for details.
Therefore, in this work, we take a first step toward evaluating multi-modal mathematical reasoning by focusing on MLLMs’ code-related capabilities in performing visual operations. We focus on two primary aspects: (1) Multi-modal Code Generation and (2) Multi-modal Code Editing. Multi-modal Code Generation evaluates models’ ability to transform the image input into the corresponding construction code. Multi-modal Code Editing encompasses three common types of visual operations: 1) Deletion tests models’ capability to identify and remove distracting elements, streamlining figures to highlight key information; 2) Modification evaluates models’ proficiency in updating images through adding auxiliary lines or altering visual elements; 3) Annotation assesses models’ ability to add numerical values, symbols, and other markings at appropriate locations, enhancing visual information expression through clear annotations.
Move one step further, we manually curate a high-quality dataset. Considering the diversity of mathematical visual reasoning tasks, we carefully select five representative types of graphics: (1) geometric figures, (2) function graphs , (3) bar charts, (4) line graphs, and (5) pie charts. For each sample, we conduct detailed manual annotations to construct initial dataset as shown in Fig. 2, including: (1) original question image; (2) images after four visual operations; (3) codes and instructions for implementing these visual operations. To ensure comprehensive and reliable evaluation, we transform the initial dataset into two question types: 10,293 multiple-choice questions to test models’ selection ability of visual operations, and 7,552 open-ended questions to assess models’ ability to generate visual operations. For free-generation questions, we also establish a evaluation strategy based on Chain-of-Thought (CoT) to conduct fine-grained assessment across multiple dimensions, including content accuracy, position consistency, code correctness and completeness.
We conduct comprehensive experiments on nine mainstream closed-source and open-source MLLMs using diverse prompting strategies. Our experimental results reveal these key findings: (1) Existing MLLMs still show significant performance gaps compared to human-level capability on our visual operation evaluation suite, indicating that our benchmark poses substantial challenges to current MLLMs; (2) Among different types of visual operations, visual modification proves to be the most challenging, while among different types of diagrams, function plots present the highest level of difficulty in processing; (3) Increasing the number of shots can moderately improve models’ understanding and execution capabilities in visual operations.
Our main contributions can be summarized as follows:
  • To our knowledge, we are the first to propose a fine-grained evaluation benchmark for visual operations, providing new perspectives and tools for in-depth assessment of MLLMs’ visual understanding capabilities.
  • We construct a carefully annotated large-scale visual operation evaluation dataset covering various mathematical graphic types and visual operation types.
  • Through extensive experiments, we reveal the capability boundary and improvement potential of MLLMs in visual operations, providing important references for future research.
II Task Formulation
This work evaluates the multi-modal visual operation capabilities of MLLMs in the form of multi-modal coding, focusing on two key aspects: multi-modal code generation and multi-modal code editing. These aspects reflect common patterns in multi-modal mathematical reasoning, where models must translate image (e.g., function graphs or charts) into code and make precise, coordinated edits to code based on image edits.
  • Multi-modal Code Generation (MCG): Given an image $V_{in}$, MLLM generate the corresponding code $C_{in}$ that can precisely construct such image. The code can either be Python Matplotlib, or . Formally, denoting the generation instruction as $I_{g}$.
    $C_{in}=MLLM(V_{in},I_{g})$ (1)
  • Multi-modal Code Editing (MCE): MLLM is given an initial image $V_{in}$, its code $C_{in}$, and an edited image $V_{out}$. MLLM is expected to generate the code for $V_{out}$, denoted as $C_{out}$ by editing the initial code $C_{in}$. Formally, denoting the editing instruction as $I_{e}$.
    $C_{out}=MLLM(V_{in},V_{out},C_{in},I_{e})$ (2)
For the MCE, we scrutinize multi-modal mathematical reasoning examples and thoroughly define three types of editing operations.
  • MCE (Deletion): $V_{out}$ is obtained by removing certain elements from $V_{in}$—for example, deleting a pie area in the pie chart, or deleting a bar in the bar chart.
  • MCE (Modification): $V_{out}$ is derived from $V_{in}$ by modifying specific elements, such as adding an auxiliary line in the geometric plot, or adjusting bar height in the bar chart.
  • MCE (Annotation): $V_{out}$ is obtained by locating and annotating the value for certain elements in $V_{in}$. Different from the other tasks, We deliberately omit $V_{out}$ from the input to assess the model’s ability to place value annotations in appropriate positions. This approach allows us to assess more deeply how well the model understands proper label placement.
In total, we define four evaluation tasks.
III Dataset Construction
A significant challenge in implementing these evaluation tasks is the lack of dataset to perform such evaluation. To address this limitation, we manually construct a dataset MathOPEval (Visual Mathematical OPeration Evaluation) that meets our specific evaluation requirements. The construction process involves two stages: (1) creating an initial dataset $D_{\text{init}}$ containing codes, images, and instructions for the four task types and (2) transforming this dataset into two formats - free-form generation $D_{\text{gen}}$ and multiple-choice questions $D_{\text{mc}}$.
III-A Initial Dataset Construction
Our initial dataset $D_{\text{init}}$ consists of images $V$, their corresponding construction codes $C$, and textual instructions $I$. For each image, we construct three image variations along with their codes and instructions corresponding to the four evaluation tasks. Specifically,
  • Original state ($V_{\text{orig}}$, $C_{\text{orig}}$, $I_{\text{orig}}$): The original image, its code, and a textual description of the image.
  • Deleted state ($V_{\text{del}}$, $C_{\text{del}}$, $I_{\text{del}}$): The image with elements deleted from $V_{\text{orig}}$, its code, and a hint on the possible deletion operations.
  • Modified state ($V_{\text{mod}}$, $C_{\text{mod}}$, $I_{\text{mod}}$): The image with elements modified from $V_{\text{orig}}$, its code, and a hint on the possible modifications.
  • Annotated state ($V_{\text{ann}}$, $C_{\text{ann}}$, $I_{\text{ann}}$): The annotated image, its code and the corresponding annotation instruction containing the elements and values for annotation, such as angle measures, coordinates, or numerical values.
To ensure the diversity of the image types, we consider three sources, five types of mathematical visualizations: geometric figures, function plots, statistical charts including bar charts, line graphs, and pie charts. For each type of mathematical visualization, we consider the diversity of their context domains. We delineate our source of data and construction details as follow.
  • Geometric Figures: Starting with geometric problems from the Geometry3K dataset [9], we first obtain $V_{\text{orig}}$ and manually annotate the image descriptions ($I_{\text{orig}}$) and code $C_{\text{orig}}$. The deletion state involves removing specified line segments to simplify the visualization, while the modification state incorporates auxiliary geometric elements such as medians, perpendicular lines, angle bisectors, and parallel lines to facilitate problem-solving. For the annotation state, we enhance figures with point labels and line measurements from problem descriptions.
  • Function Plots: We develop a standardized approach leveraging common high school-level functions collected from the internet [^1] https://homework.study.com [^2] https://mathspace.co/us. We set the domain of the functions as [5,5][-5,5], and generate initial plots and instructions ($V_{\text{orig}}$, $C_{\text{orig}},I_{\text{orig}}$). For the deletion stage, we delete part of the domain. The modification stage implements various function transformations, while the annotation phase focuses on marking critical points, including x-axis and y-axis intersections and extreme points.
  • Statistical Charts: We consider three types of charts: bar charts, line graphs, and pie charts. We collect the original state ($V_{\text{orig}}$, $C_{\text{orig}}$, $I_{\text{orig}}$) from the ChartX dataset [10]. In the deletion state, we remove selected elements from the visualizations: bars from bar charts, points at specific x-coordinates from line graphs, and areas from pie charts. In the modification state, we adjust specific visual properties—such as changing bar heights in bar charts, altering y-coordinates of selected points in line graphs, and modifying sector proportions in pie charts. In the annotation state, we add numerical labels to represent key values: bar heights in bar charts, y-coordinates of points in line graphs, and sector percentages in pie charts.
Statistic Number
Total samples 1,888
- Geometric figures 426
- Function plots 101
- Bar charts 456
- Line graphs 452
- Pie charts 453
Total questions 17,845
Total images 7,552
Multiple-choice questions 10,293
- Average question length 89.52
- Average choice length 132.12
Free-form questions 7,552
- Average question length 868.64
- Average answer length 783.77
TABLE I: Basic statistics of MathOPEval.
III-B Dataset Format Conversion
Following the definition in Eq (1) and Eq (2), we transform the initial dataset $D_{\text{init}}$ into evaluation formats. Specifically, for free-form generation dataset ($D_{\text{gen}}$),
  • MCG: $C_{\text{orig}}=MLLM(V_{\text{orig}},I_{\text{orig}})$
  • MCE (Deletion): $C_{\text{del}}=MLLM(V_{\text{orig}},V_{\text{del}},C_{\text{orig}},I_{\text{del}})$
  • MCE (Modification):
    $C_{\text{mod}}=MLLM(V_{\text{orig}},V_{\text{mod}},C_{\text{orig}},I_{\text{mod}})$
  • MCE (Annotation): $C_{\text{ann}}=MLLM(V_{\text{orig}},C_{\text{orig}},I_{\text{ann}})$
For multiple-choice dataset ($D_{\text{mc}}$), we further augment each task with three carefully designed incorrect options while maintaining similar format and length.
III-C Dataset Statistics
The statistics of MathOPEval is shown in Table I. In the initial dataset $D_{\text{init}}$, each sample contains four sets of task-specific codes, images, and instructions. After format conversion, we obtain 7,552 questions in the generation dataset $D_{\text{gen}}$ and 10,293 questions in the multiple-choice dataset $D_{\text{mc}}$, resulting in a total of 17,845 questions. Fig. 3 shows the statistics on the proportion of different types of visualization input and lists their major context domains. We can observe that the dataset covers a diverse range of domains, including “Arts and Culture”, “Social Media and the Web”, demonstrating its broad applicability across different contexts.
IV Chain-of-Thought Evaluation Strategy
After constructing the evaluation dataset, we aim to develop an effective and efficient strategy for assessing the correctness of generated code outputs. While evaluation is straightforward in the multiple-choice format—by extracting answer choices and computing accuracy—it becomes more challenging in free-form generation settings. Given the high cost of human evaluation, we leverage state-of-the-art LLMs with strong language and coding capabilities. Specifically, we propose an automated Chain-of-Thought (CoT) evaluation strategy, utilizing the DeepSeek-V3 [11] model for reliable and unbiased code evaluation. Our evaluation process comprises two key steps: key element extraction and analytical scoring.
Figure 3: Distribution of five visualization types and their major context domains.
Key Element Extraction: To ensure precise and efficient evaluation, we instruct the evaluator model to first extract and compare key elements from both the generated and reference codes, with the specific elements varying by task type. This targeted comparison enables more reliable evaluation of code correctness and task-specific MLLM coding capabilities.
  • In MCG, the extracted elements vary by visualization type: mathematical expressions for function graphs, geometric components like points, lines, and shapes for geometric figures, and data components for statistical charts. For bar charts, we focus on bar heights, category labels, bar quantity and arrangement, and structural organization. Line graphs require analysis of data points, connection styles, and point density. Pie charts are evaluated based on slice proportions, category labels, and slice ordering. We extract these elements from the generated and the reference $C_{out}$ to examine their identity.
  • In MCE (Deletion), we analyze the removed elements by comparing the generated and reference $C_{out}$ against $C_{in}$.
  • In MCE (Modification), we examine the modified elements by comparing the generated and reference $C_{out}$ against $C_{in}$.
  • In MCE (Annotation), we compare the additional annotation code segments between $C_{in}$ and $C_{out}$.
Analytical Scoring: We design a comprehensive five-level scoring rule (1–2, 3–4, 5–6, 7–8, 9–10) that evaluates outputs across multiple dimensions: content accuracy, positional consistency, code correctness, and completeness. For each visualization type and task, we define detailed evaluation criteria corresponding to each score range, as outlined in the Appendix.
Aggregating the Two Steps: To illustrate the complete evaluation pipeline, Fig. 4 shows an example of our process for the MCE (Modification) task on geometric figures. Given the complete instruction, the evaluator model first identifies key elements in the generated and reference code relative to the original code. It then conducts a comprehensive analysis across the four core dimensions. Based on the defined scoring criteria, the model assigns an overall score on the code quality.
Human Evaluation: To assess the reliability of our automated evaluation, we randomly sampled 100 model-generated scores and conducted human evaluations. Results show that 92% of the scores were considered reasonable by human annotators, indicating strong alignment between automated and manual judgments. Full details of the human evaluation criteria are provided in the Appendix.
Figure 4: Example of CoT-based evaluation strategy for MCE (Modification) in geometric figures.
V Experiments
V-A Experimental Setup
We conduct comprehensive evaluations on a diverse set of models, including proprietary models GPT-4o [12], Qwen-VL-Max [13], and open-source models Qwen2.5-VL [13], Gemma3 [14], LLaVA-NeXT [15] Series, and specialized reasoning models QVQ-72B [13]. To minimize randomness in model outputs, we set the temperature parameter to 0. For multiple-choice tasks, we investigate various prompting strategies: (1) Direct instructs models to output answers directly; (2) CoT [16] elicits step-by-step reasoning paths; (3) Descriptive Chain-of-Thought (DCoT) [17] requires models to generate descriptions before answering; and (4) Visualization-of-Thought (VoT) [18] prompts models to imagine reasoning paths with the instruction “Visualize the state after each reasoning step.” Due to resource constraints, we evaluate Direct prompting for free-form generation tasks.
MCG MCE (Deletion) MCE (Modification) MCE (Annotation)
Geo Func Bar Line Pie Geo Func Bar Line Pie Geo Func Bar Line Pie Geo Func Bar Line Pie AVG
Human 93.33 91.67 98.33 96.67 96.67 100.00 100.00 100.00 100.00 100.00 96.67 98.33 98.33 100.00 98.33 98.33 100.00 96.67 98.33 96.67 97.92
GPT-4o
Direct 71.36 51.49 50.22 73.89 40.18 33.57 44.55 67.76 48.01 80.13 55.63 43.56 56.58 62.61 49.89 44.63 73.79 45.39 37.61 64.24 54.76
CoT 68.08 51.49 64.25 80.31 49.89 49.30 44.55 97.81 68.81 96.47 77.93 45.54 71.49 76.99 46.14 39.91 71.05 55.04 53.10 78.81 64.35
DCoT 78.17 50.50 58.77 83.85 50.99 48.59 22.77 97.81 67.26 96.91 77.00 48.51 69.30 75.88 47.68 54.59 75.15 69.74 66.59 80.35 66.02
VCoT 61.74 45.54 54.82 80.53 47.68 51.88 39.60 97.37 70.35 98.45 74.88 51.49 69.74 79.65 50.33 41.59 69.83 55.48 52.88 80.79 63.73
Qwen-VL-Max
Direct 84.51 43.56 41.45 58.85 43.93 52.82 59.41 91.89 69.91 98.45 68.54 44.55 67.11 60.62 54.30 50.06 70.43 51.75 50.22 71.52 61.69
CoT 79.11 46.53 84.43 90.93 49.67 45.77 68.32 97.15 72.35 98.90 81.92 47.52 76.54 78.98 48.79 48.71 67.62 61.62 47.57 62.25 67.73
DCoT 73.00 43.56 69.52 71.02 44.37 51.64 63.37 92.11 52.65 98.23 77.70 45.54 75.44 77.21 45.92 51.55 70.17 73.25 66.37 71.96 65.73
VCoT 77.46 45.54 85.96 86.06 48.34 50.94 63.37 96.71 72.79 98.90 81.46 46.53 77.63 78.32 51.43 48.06 66.06 64.47 50.66 66.67 67.87
Qwen2.5-VL-3B-Instruct
Direct 44.55 40.59 37.50 57.08 52.54 30.28 26.73 75.00 24.34 96.47 46.95 42.57 37.50 58.85 19.65 52.39 68.33 69.74 63.05 82.34 51.32
CoT 47.52 32.67 30.04 45.35 41.72 38.21 22.77 37.06 14.38 61.59 42.02 31.68 34.65 47.35 26.49 18.76 67.39 63.60 54.42 66.89 41.23
DCoT 41.58 39.60 32.68 38.27 45.03 31.46 31.68 69.30 21.68 92.05 49.06 41.58 30.70 50.22 19.87 41.91 61.22 74.78 69.69 73.95 47.82
VCoT 42.57 30.69 33.77 46.46 47.46 31.92 23.76 70.39 21.02 93.16 47.42 37.62 31.36 52.21 21.85 20.31 73.00 72.37 64.38 78.59 47.02
Qwen2.5-VL-72B-Instruct
Direct 91.55 44.55 83.77 95.13 71.96 41.55 56.44 99.56 75.22 99.78 77.93 52.48 63.82 72.79 35.98 72.70 79.62 92.11 91.15 97.79 74.79
CoT 77.70 47.52 87.94 93.58 24.06 32.63 50.50 96.93 41.37 98.90 82.16 48.51 62.94 79.20 26.05 50.06 68.68 77.63 54.65 85.87 64.34
DCoT 85.92 41.58 83.77 94.47 25.61 32.63 59.41 92.98 44.47 99.56 81.22 46.53 58.11 79.87 35.54 54.40 75.00 78.95 63.50 90.95 66.22
VCoT 78.40 42.57 88.82 93.81 61.59 33.33 53.47 96.71 53.76 98.90 81.69 46.53 59.87 78.98 27.59 48.19 70.52 78.07 54.42 86.75 66.70
Gemma3-4B-IT
Direct 32.86 26.73 30.48 24.78 28.48 34.04 14.85 25.66 20.35 37.53 31.22 24.75 33.77 35.40 33.55 29.43 9.89 33.11 29.42 30.24 28.33
CoT 10.09 21.78 26.75 24.56 23.84 25.59 17.82 37.94 19.69 23.84 28.40 31.68 41.89 44.03 41.28 9.77 23.07 25.88 14.60 29.14 26.08
DCoT 24.65 5.94 17.76 6.42 20.53 12.21 0.99 20.39 0.88 9.71 10.80 13.86 7.24 6.86 0.22 0.19 2.83 0.88 1.11 0.44 8.20
VCoT 10.09 18.81 18.86 9.29 22.52 14.79 9.90 8.55 21.90 22.30 23.71 30.69 38.38 34.96 32.45 10.03 19.70 28.29 24.34 28.70 21.41
Gemma3-27B-IT
Direct 44.84 17.82 21.05 26.55 34.00 35.45 24.75 22.15 33.19 25.39 29.58 22.77 28.29 24.56 43.71 24.26 4.94 28.51 25.66 31.79 27.46
CoT 15.26 23.76 31.36 28.10 36.87 16.90 26.73 18.42 21.24 27.59 27.70 13.86 33.99 37.61 44.15 11.32 10.96 29.39 22.57 32.01 25.49
DCoT 26.76 15.84 27.41 32.96 35.76 7.28 11.88 6.58 12.39 5.74 19.72 21.78 8.99 10.84 11.04 8.73 8.84 17.11 18.36 22.96 16.55
VCoT 5.87 24.75 27.41 25.88 35.54 25.59 24.75 11.40 15.27 25.17 31.46 22.77 41.01 40.93 35.98 18.76 15.26 24.12 23.67 32.23 25.39
LLaVA-NeXT-8B
Direct 42.96 26.73 23.03 30.09 40.18 25.35 41.58 83.55 51.55 68.21 42.49 41.58 31.80 31.86 33.11 43.40 51.19 51.32 35.84 50.55 42.32
CoT 30.05 12.87 24.34 11.50 20.53 20.66 39.60 28.07 12.83 39.07 43.43 12.87 15.13 15.27 22.96 10.35 43.29 44.74 21.90 29.58 24.95
DCoT 26.06 12.87 21.71 12.39 10.60 17.00 24.75 41.23 14.16 23.62 29.58 0.99 16.01 12.83 20.53 13.26 28.47 44.08 25.44 25.39 21.05
VCoT 35.92 17.82 25.44 10.40 26.27 25.82 34.65 44.96 30.53 36.20 31.22 5.94 17.98 18.36 28.26 13.13 41.79 33.77 19.69 27.37 26.28
LLaVA-NeXT-72B
Direct 73.24 34.65 35.75 36.95 33.33 36.62 41.58 84.65 33.85 83.89 50.47 30.69 30.04 38.94 35.76 58.21 63.49 63.38 49.56 69.98 49.25
CoT 68.54 26.73 30.92 41.37 19.43 34.98 43.56 86.40 44.91 83.00 46.24 24.75 36.18 45.58 29.80 52.20 56.25 67.98 53.32 71.96 48.21
DCoT 56.81 15.84 33.99 38.94 32.45 27.23 54.46 87.06 41.81 84.55 44.37 24.75 29.17 32.96 26.71 32.34 65.75 72.15 56.19 77.92 46.77
VCoT 64.55 27.72 33.55 38.94 32.01 35.68 46.53 85.96 44.03 82.34 44.84 23.76 26.75 35.84 34.00 49.68 56.46 69.30 55.97 72.85 48.04
AVG 52.24 32.15 44.30 49.65 37.42 32.87 37.16 64.67 38.65 68.28 51.21 33.38 43.11 49.27 33.78 35.09 50.94 53.69 44.31 58.53 45.53
AVG 43.15 48.33 42.15 48.51 45.53
TABLE II: Main experimental results of four visual operations across five visualization types in multiple-choice format. “Geo” and “Func” are abbreviations for geometry and function. “AVG” represents the average accuracy across different models, visualization types, and visual operations. Bold numbers indicate the best results among all models and prompting methods, while underlined numbers denote the best accuracy for individual models across all prompting strategies.
MCG MCE (Deletion) MCE (Modification) MCE (Annotation) AVG
Geo Func Bar Line Pie Geo Func Bar Line Pie Geo Func Bar Line Pie Geo Func Bar Line Pie
GPT-4o 3.18 2.59 8.29 8.34 8.12 4.78 4.27 9.57 7.28 9.32 5.42 4.44 2.59 3.41 2.44 4.35 5.87 9.12 8.37 7.99 6.08
Qwen-VL-Max 2.86 2.40 8.23 8.56 8.50 3.59 4.64 9.78 7.71 9.10 4.54 3.59 2.61 4.18 2.49 5.00 5.27 9.06 7.28 8.34 5.97
Qwen2.5-VL-3B-Instruct 2.19 1.88 6.71 8.05 8.80 1.18 3.63 3.30 2.38 6.42 2.45 2.70 1.80 1.60 1.81 3.09 4.17 5.27 5.18 3.65 3.83
Qwen2.5-VL-72B-Instruct 2.76 1.94 8.45 8.83 8.60 3.73 4.05 9.59 8.33 8.75 4.70 3.29 2.48 4.21 2.57 4.15 5.55 9.17 7.73 7.75 5.94
Gemma3-4B-IT 1.90 1.08 6.37 6.67 8.49 1.21 1.01 1.39 1.07 1.97 2.00 1.00 2.01 1.96 2.32 2.51 1.01 7.00 7.39 4.58 3.30
Gemma3-27B-IT 2.62 1.27 7.06 7.52 8.79 1.92 1.00 3.31 2.91 6.07 2.09 1.00 1.99 1.97 2.45 3.31 1.00 8.61 6.98 5.40 4.05
LLaVA-NeXT-8B 1.71 1.06 5.92 6.85 8.57 1.67 3.43 1.10 1.12 1.81 1.81 3.34 1.74 1.69 1.74 2.63 3.89 5.06 5.93 8.72 3.51
LLaVA-NeXT-72B 2.34 1.21 7.36 7.17 8.79 1.54 3.38 4.34 1.88 6.65 2.30 1.67 1.83 1.66 1.91 3.62 4.00 7.86 7.71 4.06 4.09
AVG 2.44 1.68 7.30 7.75 8.58 2.45 3.18 5.30 4.09 6.26 3.16 2.63 2.13 2.58 2.22 3.58 3.85 7.64 7.07 6.31 4.51
AVG 5.55 4.26 2.54 5.69 4.59
TABLE III: Main experimental results of four visual operations across five visualization types in generation format. Bold numbers indicate the best results among all models.
V-B Main Results
Tables II and III present main results for multiple-choice questions and free-form generation tasks. Our analysis of these results reveals several key findings:
Model Performance: A substantial performance gap exists between models and human baseline across all tasks. Visual operations that humans find intuitive pose significant challenges for models, with an average performance gap of 52.39% and peaking at 67.13% for geometric deletion tasks. Among the evaluated models, GPT-4o demonstrates superior performance in free-form generation with an average score of 6.08, while Qwen2.5-VL-72B-Instruct achieves the highest accuracy of 68.02% in multiple-choice questions. At the other end of the spectrum, Gemma3-4B-IT consistently shows the lowest performance, scoring 21.00% and 3.30 in multiple-choice and free-form generation respectively. The results demonstrate that generally, visual operation comprehension improves with model size, aligning with established scaling laws in language models [19].
Task Difficulty: The complexity varies significantly across different task types. In both settings, MCE (Modification) emerges as the most challenging with an average accuracy of merely 42.15% in multiple-choice questions and an average score of just 2.54 in free-form generation. MCE (Annotation), however, achieves the highest scores in both settings, indicating that while models excel at understanding and processing numerical annotations, they struggle considerably with tasks requiring detailed comprehension and manipulation of image modifications.
Visualization Challenges: Among different visualization types, function plots present the most significant challenges, achieving only 38.41% accuracy in multiple-choice setting and an average score of 2.83 in free-form generation. And the difficulty level of geometric problems follows accordingly. While pie charts and line graphs generally demonstrate better performance with average accuracies around 50%, line graphs unexpectedly show lower multiple-choice accuracy at 45.47% compared to other statistical charts, revealing specific difficulties in understanding and manipulating line graph modifications.
Prompting Effectiveness: Contrary to previous findings, Direct prompting outperforms other strategies with 43.33% accuracy, while DCoT shows the lowest performance at 37.59%. This unexpected result likely stems from the more complex spatial relationships inherent in mathematical visualizations compared to natural images, where imprecise descriptions can actually hinder performance. Notably, while CoT prompting does not consistently outperform Direct prompting, larger models demonstrate enhanced CoT capabilities more frequently, suggesting the potential emergence of reasoning abilities with increased model scale.
V-C In-Depth Analysis
This section contains further analysis of the evaluation of visual operations.
Reasoning-Enhanced Models vs. General Models: To investigate whether reasoning-enhanced models outperform general models of similar size in visual operations, we conduct a comparative analysis between QVQ-72B-Preview and Qwen2.5-VL-72B-Instruct. Fig. 5 reveals an interesting pattern: the reasoning-enhanced model demonstrates superior performance in MCG of mathematical functions and MCE (Deletion) of line charts, but shows comparable or slightly inferior performance across other tasks. This performance disparity suggests that while reasoning enhancement effectively improves the model’s reasoning capabilities, it does not necessarily translate to better performance in general visual operations. Such observation raises important questions about the specialization of model capabilities and the trade-offs between focused reasoning enhancement and general visual understanding.
Figure 5: Comparison between reasoning-enhanced and general-purpose models on multiple-choice format using Direct prompting strategy.
Figure 6: Analysis of shot count effects on Qwen-VL-Max’s multiple-choice performance across four visual operations in geometric figures, using CoT prompting strategy.
Impact of In-Context Learning: We investigate the effectiveness of in-context learning (ICL) using Qwen-VL-Max on geometric figures. As demonstrated in Fig. 6, ICL significantly enhances performance in MCE (Annotation) and MCE (Deletion) tasks compared to zero-shot learning, which can be attributed to the relatively straightforward nature of these operations. However, when it comes to more complex tasks such as MCG and MCE (Modification), which involve diverse geometric elements (e.g., triangles, rectangles) and various line types (including parallel lines, perpendicular lines, angle bisectors, and median lines), ICL may introduce unintended noise. This complexity and the wide variability of possible operations can actually lead to performance degradation.
Model X Y Extreme AVG
GPT-4o 74.65 60.15 82.57 72.45
Qwen-VL-Max 71.28 57.67 76.76 68.57
Qwen2.5-VL-3B-Instruct 69.68 64.11 68.66 67.48
Qwen2.5-VL-72B-Instruct 74.47 53.47 92.43 73.45
Gemma3-4B-IT 10.28 14.60 16.73 13.87
Gemma3-27B-IT 2.84 12.38 14.79 10.00
LLaVA-NeXT-8B 54.79 33.91 34.86 41.19
LLaVA-NeXT-72B 65.43 60.40 55.63 60.49
AVG 52.93 44.59 55.30 50.94
TABLE IV: Comparison of x-intercepts, y-intercepts, and extreme points in function-type MCE (Annotation) for multiple-choice.
Model Backward Forward AVG
GPT-4o 62.82 45.18 54.00
Qwen-VL-Max 72.48 49.60 61.04
Qwen2.5-VL-3B-Instruct 57.81 33.34 45.58
Qwen2.5-VL-72B-Instruct 79.84 56.34 68.09
Gemma3-4B-IT 17.60 12.35 14.98
Gemma3-27B-IT 19.21 15.77 17.49
LLaVA-NeXT-8B 27.93 20.04 23.98
LLaVA-NeXT-72B 49.94 48.11 49.03
AVG 48.45 35.09 41.77
TABLE V: Comparison of forward and backward problems in geometric MCE (Annotation) for multiple-choice.
Model Angle Line AVG
GPT-4o 5.41 3.28 4.35
Qwen-VL-Max 5.19 4.81 5.00
Qwen2.5-VL-3B-Instruct 3.31 2.87 3.09
Qwen2.5-VL-72B-Instruct 5.42 2.87 4.15
Gemma3-4B-IT 2.62 2.39 2.51
Gemma3-27B-IT 3.97 2.64 3.31
LLaVA-NeXT-8B 2.69 2.56 2.63
LLaVA-NeXT-72B 4.48 2.76 3.62
AVG 4.14 3.02 3.58
TABLE VI: Comparison of angle and line annotation in geometric MCE (Annotation) for free-form generation.
Shot Count Analysis: Fig. 6 demonstrates that increasing the number of shots does not monotonically improve performance. Performance typically peaks at 2 shots and declines beyond 3 shots. This pattern suggests that careful selection of shot count is crucial for optimal performance. The performance degradation with higher shot counts may be attributed to two factors. On one hand, while initial shots enhance task understanding, excessive examples introduce confusion, especially with diverse configurations. On the other hand, models’ limited context window constrains their ability to process numerous shots effectively, leading to performance degradation.
Intersection Points vs. Extreme Points: Our analysis of MCE (Annotation) for function plots in Table VI reveals varying performance patterns for different mathematical features. Generally, models perform best at identifying extreme points, followed by x-intercepts, with y-intercepts proving most challenging. Interestingly, we observe that Qwen2.5-VL-3B-Instruct outperforms Qwen2.5-VL-72B-Instruct on y-intercept tasks, and Gemma3-4B-IT surpasses Gemma3-27B-IT on both types of intercepts. This counter-intuitive finding suggests that model size does not necessarily correlate with improved understanding of coordinate intersections.
Forward vs. Backward Questions: In MCE (Annotation) for geometric figures, we compare performance on forward and backward questions. The former identifies annotations based on geometric elements, e.g., what is the measure of angle ABC, while the latter identifies geometric elements based on annotations, e.g., which angle measures 120 degrees. As shown in Table VI, models consistently perform better on backward questions. We hypothesize this asymmetry might stem from the fact that backward questions provide explicit numerical anchors that help constrain the search space, while forward questions require more comprehensive geometric understanding and spatial reasoning.
Points vs. Lines: Regarding the generation setting for MCE (Annotation) in geometric problems, we further categorize the annotations into two types: angle markings and line labels. As illustrated in Table VI, angle annotations demonstrate notably higher scores compared to line annotations. We attribute this performance difference to the varying complexity of placement requirements: angle markings only need to be positioned near the identified vertex, whereas line labels require to be placed at the midpoint of the line segment. This more stringent positioning requirement for line labels contributes to the increased difficulty in their accurate placement.
V-D Error Analysis
We conduct a manual analysis on 100 randomly selected outputs containing reasoning processes, equally divided between 50 correct and 50 incorrect processes. Our examination reveals concerning patterns in the models’ performance in Fig. 7: among the correct answers, only 33% demonstrates both accurate results and valid reasoning processes, while 17% achieves correct answers through wrong reasoning - indicating that correct outputs do not necessarily reflect sound problem-solving capabilities. We also identify four distinct categories of failures, with visual perception errors emerging as the predominant issue, accounting for 86% of all errors. This remarkably high percentage highlights a crucial limitation in the models’ ability to accurately process and interpret visual elements. Both instruction comprehension errors and output format violations each contribute to 6% of the failures, while reasoning process errors constitute the remaining 2%. This distribution of errors suggests that visual perception mechanisms require substantial improvement for reliable visual problem-solving.
Figure 7: Error analysis of answer and reasoning precess.
Figure 8: Case study for MCG of function plots.
V-E Case Study
As shown in Fig. 8, Qwen2.5-VL-3B-Instruct is presented with an image of a function and asked to provide its mathematical formula in LaTeX format. The correct function, as defined by the reference formula, is $f(x)=x^{2}-1$, which represents a standard parabola shifted one unit down on the y-axis. However, the model under evaluation provides the answer $f(x)=x^{2}$. The evaluation system assigns this response a score of 5 out of 10, classifying it as having the same function type but with a “moderate structural difference”. This case clearly demonstrates the core deficiency identified in our work: a bottleneck in fine-grained visual perception. While the model correctly identifies the general function type as a quadratic parabola ($x^{2}$), showing its capability in high-level visual classification, it fails to perceive the precise details of the graph’s position. Specifically, it does not recognize that the vertex is located at (0, -1) on the coordinate plane, instead defaulting to the most basic representation $f(x)=x^{2}$ with a vertex at the origin. This failure is not one of mathematical reasoning but of visual perception - the model does not correctly identify the vertical shift from the y-axis, as evidenced by the missing 1-1 constant in its output. This case exemplifies how models often grasp the general “what” of an image but struggle with the specific “where” and “how,” which are crucial for accurate mathematical interpretation.
VI related work
Multimodal Mathematical Assessment: With the rapid development of MLLMs, numerous benchmarks have emerged to evaluate visual mathematical reasoning capabilities[20, 21, 22, 23, 20, 24, 25, 26, 27, 28]. Previous work [29, 30, 31], such as Inter-gps [9] and GeoQA [32], focuses on geometric problems, while others [33, 34] cover broader tasks including function graphs and statistical charts. Though recent work [5, 7, 8] introduces a “multi-modal input-multi-modal output” paradigm, existing evaluations still assess only final answer accuracy, overlooking the quality of intermediate visual operations during reasoning.
Visual Programming: The decomposition of complex visual tasks into manageable steps through programming interfaces has demonstrated significant potential in advancing visual reasoning capabilities[35, 36, 37, 38]. The core idea behind this approach is to leverage the logical and precise nature of computer programs to transform complex visual problems, which traditionally rely on human intuition and experience, into a series of clear, step-by-step operations. Among them, Visprog [39] and ViperGPT [40] showcase the effectiveness of using MLLMs to generate Python code for sequential visual operations. SKETCHPAD [5] further enables dynamic visual manipulations based on intermediate results, achieving particular success in geometric reasoning through visual operations like drawing auxiliary lines via code. This progress inspires our work to move beyond evaluating just final mathematical solutions and toward assessing the quality and appropriateness of intermediate visual operations in the reasoning process.
VII Future Directions
Enhancing core visual perception and grounding: Our findings show that visual perception is the primary bottleneck, causing the vast majority of errors. Future efforts should be directed towards substantially improving the foundational visual architectures of MLLMs, with a specific emphasis on refining their capabilities for fine-grained feature extraction and precise semantic grounding.
Improving reasoning paradigms for visual tasks: We find that complex prompting strategies like CoT don’t consistently outperform simpler approaches in visual contexts. This suggests a need for reasoning paradigms better suited to visual tasks. Future work could explore strategies that more tightly integrate reasoning steps with visual manipulation, moving beyond purely text-based approaches.
Optimizing visual in-context learning: Our research shows that simply increasing the number of shots doesn’t always improve performance. Future work should investigate more efficient ICL strategies, including methods for selecting the most relevant examples and techniques for compressing visual information to better utilize the model’s context window.
VIII conclusion
In this paper, we presented the first fine-grained evaluation benchmark for visual operations in mathematical reasoning tasks. Our framework comprehensively assesses four fundamental visual operations across five types of mathematical visualizations, providing a systematic approach to evaluating MLLMs’ capabilities. Through extensive experiments, we identified significant performance gaps between current models and human, particularly in function plots and MCE (Modification) task. Additionally, we observed that while in-context learning can enhance performance, its effectiveness varies across different tasks and requires careful consideration of shot count. We hope our framework and findings will facilitate the development of more capable models in visual mathematical reasoning.